OpenStudy (abbycross167):

Can someone please try to explain to me how to answer this question? Please Question~ Solve the systems using the elimination method. x + y = 0 3x + y = -4

OpenStudy (anonymous):

subtract the first equation from the second, that will eliminate the y term

OpenStudy (anonymous):

\[3x+y=-4\\ x+y=0\] what do you get when you subtract line 2 from line 1?

OpenStudy (abbycross167):

3x=-4?

OpenStudy (abbycross167):

@satellite73 ??

OpenStudy (anonymous):

close but no the y terms go, and you do get \(-4\) on the right but what is \(3x-x\)?

OpenStudy (abbycross167):

2x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so you have \[2x=-4\] which should be easy to solve

OpenStudy (abbycross167):

x=-2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and since \(x+y=0\) and \(x=-2\) that means \[-2+y=0\] and you can solve that for \(y\)

OpenStudy (abbycross167):

y=2?

OpenStudy (anonymous):

yes

OpenStudy (abbycross167):

ok then what?

OpenStudy (anonymous):

have a beer?

OpenStudy (anonymous):

or a snack, if you are under age you are done \[x=-2,y=2\]

OpenStudy (abbycross167):

haha, I am underage. so the answer is just x=-2, y=2? I thought when using the elimination method you'll get an answer that looks like this~~ (11/7, 12/7) the numbers could be different however.

OpenStudy (anonymous):

if you want your answer as an ordered pair, feel free to write \(x=2,y=2\) as the pair \((-2,2)\)

OpenStudy (abbycross167):

oh ok... Thank you so much! You are a great help!

OpenStudy (anonymous):

yw hope it helped you got more?

OpenStudy (abbycross167):

yes sir, a few more but I'm going to try to answer them on my own.

OpenStudy (anonymous):

ok good post if you get stuck

OpenStudy (anonymous):

sir lol i like that